# Does Less Hallucination Mean Less Creativity? An Empirical Investigation in LLMs

### v1

@@ 1. Introduction
The development of Large Language Models (LLMs) is a landmark achievement in the field of Natural Language Processing (NLP). They exhibit unprecedented abilities in natural language understanding (hendrycksMeasuringMassiveMultitask2021) and reasoning (yufeiNaturalLanguageReasoning2024; kojimaLargeLanguageModels2023; zhao2025surveylargelanguagemodels). Unfortunately, modern LLMs often suffer from hallucination, the tendency to generate factually incorrect content (Huang_2025; rawteTroublingEmergenceHallucination2023). As such, there has been significant effort in the field to understand (yaoLLMLiesHallucinations2024; kalaiWhyLanguageModels2025) and combat (gumaan2025theoreticalfoundationsmitigationhallucination) hallucination, especially in high-stakes domains such as AI-assisted scientific discovery where factual reliability is essential (zhang2025advancingscientificmethodlarge).
@@
However, little is known about how interventions that suppress hallucination affect a model’s creative potential — a crucial ingredient for hypothesis generation and scientific ideation. This question matters because creativity often involves making unconventional connections. For example, scientists sometimes generate useful hypotheses by linking concepts that at first seem unrelated; a process that, in models, can resemble the kind of associative leaps that might otherwise be labeled as hallucination. We adopt the definition of creativity from human psychology (guilfordCreativity1950), where creativity is divided into convergent thinking: solving the problem correctly and within means, and divergent thinking: generation of different ideas. In this paper, we aim to investigate the relationship between hallucination-reduction methods and creativity, shedding light on how factual control interacts with creative reasoning.
@@
Our experimental setup is illustrated in Figure 1. We evaluate creativity using two benchmarks from distinct domains: (1) NeoCoder (lu2025benchmarkinglanguagemodelcreativity), and (2) CS4 (atmakuru2024cs4measuringcreativitylarge). NeoCoder evaluates creativity in solving increasingly constrained programming problems: a rule-based setting similar to scientific experimentation under fixed laws. Meanwhile, CS4 tests open-ended story generation, reflecting the imaginative thinking needed for hypothesis generation in science.
@@
We re-implemented three hallucination-reduction techniques: Chain of Verification (CoVe) (dhuliawala2023chainofverificationreduceshallucinationlarge), Decoding by Contrasting Layers (DoLa) (chuang2024doladecodingcontrastinglayers), and Retrieval-Augmented Generation (RAG) (lewis2021retrievalaugmentedgenerationknowledgeintensivenlp). For each, we measure both convergent and divergent creativity before and after applying the method.
@@
Before conducting our experiments, we hypothesized that hallucination-reduction techniques would generally suppress a model’s creative abilities, given that creative thinking often relies on making unconventional associations that may be mistaken for hallucination. However, our empirical results reveal a surprising pattern. Different hallucination-reduction methods affect divergent creativity in opposite ways. CoVe enhances the model’s ability to generate diverse and original ideas, whereas DoLa suppresses it. Meanwhile, convergent creativity is largely unaffected. We further examine whether this trend generalizes across different model families: LLaMA (grattafiori2024llama3herdmodels), Qwen (hui2024qwen25codertechnicalreport), and Mistral (jiang2023mistral7b), and across model scales: 1B, 8B, and 70B parameters. The consistent trend suggests that the creativity-hallucination relationship is an inherent characteristic of LLMs, not an artifact of model size or architecture.
@@ 2. Related Works and Background
(dhuliawala2023chainofverificationreduceshallucinationlarge) introduces Chain of Verification (CoVe), a structured approach that enhances factual consistency through multi-stage reasoning. The process comprises four stages: (1) drafting an initial answer, (2) generating verification questions based on the draft, (3) answering these questions and synthesizing recommendations, and (4) producing a refined final response. This iterative verification chain enables models to critically evaluate and correct their own outputs before finalizing an answer.
@@
(chuang2024doladecodingcontrastinglayers) proposes Decoding by Contrasting Layers (DoLa), a simple decoding method to make large language models more factual. Instead of using only the final layer’s output, DoLa contrasts the predictions from a higher layer with those from an earlier one. The earlier “premature” layer is chosen dynamically at each step by finding which layer’s output differs most from the final layer using Jensen-Shannon divergence (61115). The model then subtracts the earlier layer’s logits from the later layer’s, emphasizing tokens that are learned throughout the layers, while reducing less-reliable tokens from the lower layers.
@@
Another widely adopted approach is Retrieval-Augmented Generation (RAG) (lewis2021retrievalaugmentedgenerationknowledgeintensivenlp), which enhances factual accuracy by retrieving relevant external information from a knowledge source before generating the final response. Integrating retrieved evidence into the model’s context, RAG enables the system to ground its outputs in verifiable data rather than relying solely on parametric memory.
@@
This dataset (lu2025benchmarkinglanguagemodelcreativity) was introduced to evaluate the creativity of LLMs in a structured, constraint-driven programming setting. It comprises 199 CodeForces problems, each paired with around 30 human-written correct solutions. Every problem $x_{i}$ is associated with a progressively growing set of constraints $\mathcal{C}_{t}^{i}=\{\tau_{1}^{i},\tau_{2}^{i},\ldots,\tau_{t}^{i}\}$, where $t$ denotes the constraint state ($t=|\mathcal{C}_{t}^{i}|$), and the maximum number of constraints is $T=5$. Each instance at state $t$ is represented as
@@
The CS4 benchmark (atmakuru2024cs4measuringcreativitylarge) provides a controlled framework for evaluating LLMs on creative story generation under progressively more complex constraints. It employs a constraint-generation strategy that produces stylistic and open-ended constraints from user instructions. Each of the 50 instructions is expanded to 39 cumulative constraints, segmented into sets of 7, 15, 23, 31, and 39. This results in 250 unique prompts $(50\times 5)$.
@@ 3. Effects of Hallucination-Reduction Methods on Creativity
In this section, we provide details on the structure of our experiment (Figure 1). We evaluated models on the two datasets — NeoCoder (lu2025benchmarkinglanguagemodelcreativity) and CS4 (atmakuru2024cs4measuringcreativitylarge) — to assess both factual consistency and creative robustness. For the NeoCoder benchmark, we evaluated five models: LLaMA 70B, LLaMA 8B, LLaMA 1B (grattafiori2024llama3herdmodels), Mistral 7B (jiang2023mistral7b), and Qwen-Coder 7B (hui2024qwen25codertechnicalreport) to ensure comprehensive coverage across diverse model families and parameter scales. For the CS4 benchmark, experiments were conducted with LLaMA 8B, LLaMA 1B, and Mistral 7B under identical experimental settings for consistency. All generations were performed three times independently using identical configurations, and the mean performance across runs was reported for reliability. Wherever LLM-as-a-Judge evaluation was required, GPT-5-mini (openai2025gpt5) was used for consistency and neutrality.
We re-implemented all three hallucination-reduction methods using the following settings. CoVe was implemented using the AutoGen multi-agent framework (wu2023autogenenablingnextgenllm), while for DoLa, we chose to only contrast with even-indexed layers, as per the original implementation. Finally, for RAG, we employed ColBERTv2 (santhanam2022colbertv2effectiveefficientretrieval), following the RAGLAB Framework (zhangRAGLABModularResearchOriented2024), as the retrieval backbone. It indexed a large corpus of coding tutorials, library documentations, Github repositories, and programming solutions, taken from the CodeRAG-bench (wangCodeRAGBenchCanRetrieval2025) retrieval documents. Documents were vectorized into ColBERTv2 embeddings and retrieved via cosine similarity, with the top 3 ranked segments appended to the prompt for context-grounded generation. RAG was applied only to the NeoCoder benchmark, as open-ended story generation in CS4 lacks a defined retrieval corpus. Further fine-grained details are available in Appendix E.
@@ Results
Our observation supports the hypothesis that questioning improves creativity (wroblewskaApplyingTextMining2025), by encouraging a broader exploration of the solution space in the model. This aligns with recent findings that exploration of different reasoning paths can help models sidestep ’tunnel vision’ (wen2025parathinker) and explore better, more unique solutions to problems. In human cognition, similar mechanisms are well-documented: questioning strategies reduce fixation and enhance creative output (razRoleAskingMore2025), while brainstorming techniques stimulate divergent thinking (ritter2017enhancementcreativethinking). The CoVe verification process may function analogously, prompting the model to reconsider and explore alternative solutions rather than committing prematurely to a single response path.
@@
This neutral effect could stem from retrieval quality issues. Studies show that irrelevant retrieved documents introduce noise that misleads LLM generation (shi2023large), showing that performance could degrade when context lacks direct relevance to the query. Recent work has also shown that redundant knowledge in RAG corpora can hurt performance on questions the LLM can already answer (luo2024zerorag). In our setup, CodeForces problems are presented through narrative scenarios, while the retrieval corpus contains technical coding tutorials and documentation from CodeRAG-bench. This semantic mismatch between anecdotal problem descriptions and tutorial-style documentation results in retrieved documents that lack the specific algorithmic insights needed for competitive programming tasks. However, this net-neutral effect suggests potential for improvement: if RAG were to retrieve documents containing newer patterns or problem-solving strategies outside the model’s training distribution, it could enhance both convergent creativity (by providing relevant factual guidance) and divergent creativity (by exposing the model to unfamiliar approaches). The key lies in ensuring retrieved documents offer genuinely new and relevant information rather than redundant or misaligned content.
@@
We investigated the influence of early layers on the model’s divergent creativity. Using probing methods inspired by Inference-Time Intervention (ITI) (liInferenceTimeInterventionEliciting2024), we used linear probes to identify which attention head are most correlated with creativity. Specifically, we trained linear probes to predict whether the model is going to generate a divergently creative output on the NeoCoder dataset. Since DoLa operates on entire layers rather than individual attention heads, we aggregated the attention head-level correlations within each layer to obtain layer-level correlation scores for direct compatibility. Figure 4 shows that early layers exhibit stronger correlations with creativity than later layers. This finding supports our hypothesis that DoLa’s contrastive decoding mechanism, which specifically contrasts against early layer representations, inadvertently suppresses divergent creativity by removing the very layer activations responsible for creative generation.
@@ Acknowledgments
Figure 6: Impact of decoding methods on convergent creativity. The plots show the percentage improvement over baseline performance for various language models across six constraints. Top: NeoCoder dataset. Bottom: CS4 dataset. The horizontal line at y=0 represents the baseline (generation without hallucination-reduction methods). Positive values indicate improvement over baseline , while negative values indicate degradation.
Figure 7: Model-specific nature of creative probes. Divergent creativity improvement across constraint states for Qwen-coder 7B using Creative DoLa with probes trained on different generations. When probes are trained on LLaMA 8B creative generations (blue line), they fail to improve Qwen-coder’s divergent creativity. However, when probes are trained on Qwen-coder’s own creative generations (brown line), the method successfully enhances divergent creativity. This demonstrates that creative probes must be trained on model-specific activations to be effective.
$\textsc{Increasing divergent creativity with probes}(x_{t}|x_{<t})=\text{softmax}(\mathcal{F}(x_{t},N,M,\mathcal{A},\mathcal{B})),\quad\text{where}$ (8)
$\mathcal{F}(x_{t},N,M,\mathcal{A},\mathcal{B})=\begin{cases}\overbrace{\log\left(\frac{q_{N}(x_{t})}{q_{M}(x_{t})}\right)}^{\text{Normal DoLa}}+\alpha\left(\sum_{a\in\mathcal{A}}\gamma_{a}\log\left(q_{a}(x_{t})\right)-\sum_{b\in\mathcal{B}}\gamma_{b}\log\left(q_{b}(x_{t})\right)\right),&\text{if }x_{t}\in\mathcal{V}_{\text{head}}(x_{t}|x_{<t})\\ -\infty,&\text{otherwise.}\end{cases}$ (9)
@@
Following the success of linear probes in identifying properties such as truthfulness (liInferenceTimeInterventionEliciting2024) and even safety-related concepts (teo2025blessingcursedimensionalitysafety), we similarly hypothesize that creativity-related features may be linearly separable in the representation space.
@@
We follow similar notation as the original DoLa paper (chuang2024doladecodingcontrastinglayers).
@@
Following Contrasting Decoding method (li2023contrastivedecodingopenendedtext), the subset $\mathcal{V}_{\text{head}}(x|x_{t})\in\mathcal{X}$ contains tokens with sufficiently ’high’ probabilities at the last layer $N$. The parameter $\beta$ corresponds to the number of tokens considered. In the paper, $\beta$ is set to $0.1$.

### v2

@@ 1. Introduction
The development of Large Language Models (LLMs) is a landmark achievement in the field of Natural Language Processing (NLP). They exhibit unprecedented abilities in natural language understanding (Hendrycks et al. 2021) and reasoning (YuFei et al. 2024; Kojima et al. 2023; Zhao et al. 2025). Unfortunately, modern LLMs often suffer from hallucination, the tendency to generate factually incorrect content (Huang et al. 2025; Rawte et al. 2023). As such, there has been significant effort in the field to understand (Yao et al. 2024; Kalai et al. 2025) and combat (Gumaan 2025) hallucination, especially in high-stakes domains such as AI-assisted scientific discovery where factual reliability is essential (Zhang et al. 2025).
@@
However, little is known about how interventions that suppress hallucination affect a model’s creative potential — a crucial ingredient for hypothesis generation and scientific ideation. This question matters because creativity often involves making unconventional connections. For example, scientists sometimes generate useful hypotheses by linking concepts that at first seem unrelated; a process that, in models, can resemble the kind of associative leaps that might otherwise be labeled as hallucination. We adopt the definition of creativity from human psychology (Guilford 1950), where creativity is divided into convergent thinking: solving the problem correctly and within means, and divergent thinking: generation of different ideas. In this paper, we aim to investigate the relationship between hallucination-reduction methods and creativity, shedding light on how factual control interacts with creative reasoning.
@@
Our experimental setup is illustrated in Figure 1. We evaluate creativity using two benchmarks from distinct domains: (1) NeoCoder (Lu et al. 2025), and (2) CS4 (Atmakuru et al. 2024). NeoCoder evaluates creativity in solving increasingly constrained programming problems: a rule-based setting similar to scientific experimentation under fixed laws. Meanwhile, CS4 tests open-ended story generation, reflecting the imaginative thinking needed for hypothesis generation in science.
@@
We re-implemented three hallucination-reduction techniques: Chain of Verification (CoVe) (Dhuliawala et al. 2023), Decoding by Contrasting Layers (DoLa) (Chuang et al. 2024), and Retrieval-Augmented Generation (RAG) (Lewis et al. 2021). For each, we measure both convergent and divergent creativity before and after applying the method.
@@
Before conducting our experiments, we hypothesized that hallucination-reduction techniques would generally suppress a model’s creative abilities, given that creative thinking often relies on making unconventional associations that may be mistaken for hallucination. However, our empirical results reveal a surprising pattern. Different hallucination-reduction methods affect divergent creativity in opposite ways. CoVe enhances the model’s ability to generate diverse and original ideas, whereas DoLa suppresses it. Meanwhile, convergent creativity is largely unaffected. We further examine whether this trend generalizes across different model families: LLaMA (Grattafiori et al. 2024), Qwen (Hui et al. 2024), and Mistral (Jiang et al. 2023), and across model scales: 1B, 8B, and 70B parameters. The consistent trend suggests that the creativity-hallucination relationship is an inherent characteristic of LLMs, not an artifact of model size or architecture.
@@ 2. Related Works and Background
(Dhuliawala et al. 2023) introduces Chain of Verification (CoVe), a structured approach that enhances factual consistency through multi-stage reasoning. The process comprises four stages: (1) drafting an initial answer, (2) generating verification questions based on the draft, (3) answering these questions and synthesizing recommendations, and (4) producing a refined final response. This iterative verification chain enables models to critically evaluate and correct their own outputs before finalizing an answer.
@@
(Chuang et al. 2024) proposes Decoding by Contrasting Layers (DoLa), a simple decoding method to make large language models more factual. Instead of using only the final layer’s output, DoLa contrasts the predictions from a higher layer with those from an earlier one. The earlier “premature” layer is chosen dynamically at each step by finding which layer’s output differs most from the final layer using Jensen-Shannon divergence (Lin 1991). The model then subtracts the earlier layer’s logits from the later layer’s, emphasizing tokens that are learned throughout the layers, while reducing less-reliable tokens from the lower layers.
@@
Another widely adopted approach is Retrieval-Augmented Generation (RAG) (Lewis et al. 2021), which enhances factual accuracy by retrieving relevant external information from a knowledge source before generating the final response. Integrating retrieved evidence into the model’s context, RAG enables the system to ground its outputs in verifiable data rather than relying solely on parametric memory.
@@
This dataset (Lu et al. 2025) was introduced to evaluate the creativity of LLMs in a structured, constraint-driven programming setting. It comprises 199 CodeForces problems, each paired with around 30 human-written correct solutions. Every problem $x_{i}$ is associated with a progressively growing set of constraints $\mathcal{C}_{t}^{i}=\{\tau_{1}^{i},\tau_{2}^{i},\ldots,\tau_{t}^{i}\}$, where $t$ denotes the constraint state ($t=|\mathcal{C}_{t}^{i}|$), and the maximum number of constraints is $T=5$. Each instance at state $t$ is represented as
@@
The CS4 benchmark (Atmakuru et al. 2024) provides a controlled framework for evaluating LLMs on creative story generation under progressively more complex constraints. It employs a constraint-generation strategy that produces stylistic and open-ended constraints from user instructions. Each of the 50 instructions is expanded to 39 cumulative constraints, segmented into sets of 7, 15, 23, 31, and 39. This results in 250 unique prompts $(50\times 5)$.
@@ 3. Effects of Hallucination-Reduction Methods on Creativity
In this section, we provide details on the structure of our experiment (Figure 1). We evaluated models on the two datasets — NeoCoder (Lu et al. 2025) and CS4 (Atmakuru et al. 2024) — to assess both factual consistency and creative robustness. For the NeoCoder benchmark, we evaluated five models: LLaMA 70B, LLaMA 8B, LLaMA 1B (Grattafiori et al. 2024), Mistral 7B (Jiang et al. 2023), and Qwen-Coder 7B (Hui et al. 2024) to ensure comprehensive coverage across diverse model families and parameter scales. For the CS4 benchmark, experiments were conducted with LLaMA 8B, LLaMA 1B, and Mistral 7B under identical experimental settings for consistency. All generations were performed three times independently using identical configurations, and the mean performance across runs was reported for reliability. Wherever LLM-as-a-Judge evaluation was required, GPT-5-mini (OpenAI 2025) was used for consistency and neutrality.
We re-implemented all three hallucination-reduction methods using the following settings. CoVe was implemented using the AutoGen multi-agent framework (Wu et al. 2023), while for DoLa, we chose to only contrast with even-indexed layers, as per the original implementation. Finally, for RAG, we employed ColBERTv2 (Santhanam et al. 2022), following the RAGLAB Framework (Zhang et al. 2024), as the retrieval backbone. It indexed a large corpus of coding tutorials, library documentations, Github repositories, and programming solutions, taken from the CodeRAG-bench (Wang et al. 2025) retrieval documents. Documents were vectorized into ColBERTv2 embeddings and retrieved via cosine similarity, with the top 3 ranked segments appended to the prompt for context-grounded generation. RAG was applied only to the NeoCoder benchmark, as open-ended story generation in CS4 lacks a defined retrieval corpus. Further fine-grained details are available in Appendix E.
@@ Results
Our observation supports the hypothesis that questioning improves creativity (Wróblewska et al. 2025), by encouraging a broader exploration of the solution space in the model. This aligns with recent findings that exploration of different reasoning paths can help models sidestep ’tunnel vision’ (Wen et al. 2025) and explore better, more unique solutions to problems. In human cognition, similar mechanisms are well-documented: questioning strategies reduce fixation and enhance creative output (Raz et al. 2025), while brainstorming techniques stimulate divergent thinking (Ritter and Mostert 2017). The CoVe verification process may function analogously, prompting the model to reconsider and explore alternative solutions rather than committing prematurely to a single response path.
@@
This neutral effect could stem from retrieval quality issues. Studies show that irrelevant retrieved documents introduce noise that misleads LLM generation (Shi et al. 2023), showing that performance could degrade when context lacks direct relevance to the query. Recent work has also shown that redundant knowledge in RAG corpora can hurt performance on questions the LLM can already answer (Luo et al. 2024). In our setup, CodeForces problems are presented through narrative scenarios, while the retrieval corpus contains technical coding tutorials and documentation from CodeRAG-bench. This semantic mismatch between anecdotal problem descriptions and tutorial-style documentation results in retrieved documents that lack the specific algorithmic insights needed for competitive programming tasks. However, this net-neutral effect suggests potential for improvement: if RAG were to retrieve documents containing newer patterns or problem-solving strategies outside the model’s training distribution, it could enhance both convergent creativity (by providing relevant factual guidance) and divergent creativity (by exposing the model to unfamiliar approaches). The key lies in ensuring retrieved documents offer genuinely new and relevant information rather than redundant or misaligned content.
@@
We investigated the influence of early layers on the model’s divergent creativity. Using probing methods inspired by Inference-Time Intervention (ITI) (Li et al. 2024), we used linear probes to identify which attention head are most correlated with creativity. Specifically, we trained linear probes to predict whether the model is going to generate a divergently creative output on the NeoCoder dataset. Since DoLa operates on entire layers rather than individual attention heads, we aggregated the attention head-level correlations within each layer to obtain layer-level correlation scores for direct compatibility. Figure 4 shows that early layers exhibit stronger correlations with creativity than later layers. This finding supports our hypothesis that DoLa’s contrastive decoding mechanism, which specifically contrasts against early layer representations, inadvertently suppresses divergent creativity by removing the very layer activations responsible for creative generation.
@@ Acknowledgments
Following the success of linear probes in identifying properties such as truthfulness (Li et al. 2024) and even safety-related concepts (Teo et al. 2025), we similarly hypothesize that creativity-related features may be linearly separable in the representation space.
@@
We follow similar notation as the original DoLa paper (Chuang et al. 2024).
@@
Following Contrasting Decoding method (Li et al. 2023), the subset $\mathcal{V}_{\text{head}}(x|x_{t})\in\mathcal{X}$ contains tokens with sufficiently ’high’ probabilities at the last layer $N$. The parameter $\beta$ corresponds to the number of tokens considered. In the paper, $\beta$ is set to $0.1$.